OpenStudy (anonymous):

x/x+7 - 7/x-7 = x^2+49/x^2-49

OpenStudy (anonymous):

My first step would be to find a common denominator for those fractions so I could add them. What would that look like?

OpenStudy (anonymous):

(x+7) (x-7)?

OpenStudy (anonymous):

That's the common denominator, yep! So how can you use that to add the fractions?

OpenStudy (anonymous):

u use them as the base?

OpenStudy (anonymous):

Yep, so for example, the first fraction will go from: \[\frac{x}{x+7} \Rightarrow \frac{?}{(x+7)(x-7)}\]What goes in place of the ?

OpenStudy (anonymous):

will it be (x+7) +x-7 after factoring x^2+49

OpenStudy (anonymous):

Unfortunately, x^2+49 cannot be factored.

OpenStudy (anonymous):

What goes in place of that ? from before?

OpenStudy (anonymous):

x-7

OpenStudy (anonymous):

Close! Don't forget that there was an x there already.

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

\[\frac{x}{x+7}=\frac{x}{x+7}\times\frac{x-7}{x-7}\]

OpenStudy (anonymous):

so will that be x^2-7x/(x+7) (x-7)

OpenStudy (anonymous):

x/x+7 - 7/x-7 = x (x-7) - 7 (x + 7) / (x^2 - 49 ) = x^2 - 14x + 49 / (x^2 - 49 ) Thus [x^2 - 14x + 49 / (x^2 - 49 ) ]= [(x^2 + 49) /(x^2 - 49 )] -> -14x / (x^2 - 49 )

OpenStudy (anonymous):

so is the answer -14x/x^2-49

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thx so much

OpenStudy (anonymous):

I'm glad I can help!

OpenStudy (anonymous):

Once you have that, you can set it equal to the other side of the equation, and solve for x.